# Neural Networks Learn Generic Multi-Index Models Near Information-Theoretic Limit

### v1

@@ 1 Introduction
Modern large language models powered by deep neural networks [KSH12, HZRS16, VSP+23] have shown strong empirical success across many domains. It is believed that much of the effectiveness of the pretraining–finetuning paradigm comes from their ability to extract useful representations from data. In practice, neural networks often first capture important hidden features, which helps them to learn complex tasks more efficiently. This adaptability offers a clear performance advantage over conventional fixed feature methods, such as kernel techniques [WLLM20, AZL20, BL20].
@@
Despite these empirical successes, the theoretical understanding of why neural networks can efficiently and adaptively learn important features from data remains largely incomplete. Significant effort has been devoted in this direction, with a series of works [ABUZ18, DLS22, BES+22, GTM+23, BMZ24, LOSW24, MU25] investigating how neural networks can uncover hidden features and subsequently learn single-index and multi-index models, where the labels depend on the $d$-dimensional data inputs only through their projection onto a one-dimensional or low-dimensional subspace, respectively. The study of learning single-index models via neural networks is relatively well developed. Notably, [LOSW24] show that neural networks can successfully learn single-index models with $\widetilde{\mathcal{O}}(d)$ sample complexity, which is optimal up to leading order, while [MU25] provides a sharp DMFT theory analyzing the learning dynamics of single-index models using a two-layer neural network when the sample size and input dimension are proportional.
@@
However, a single-index model involves only one hidden feature, whereas in practice we often expect multiple distinct features, motivating the study of neural networks learning of multi-index models. On the contrary, positive results on neural networks learning of multi-index models remain rather limited. Note first that the information-theoretic limit for learning multi-index models is still $\Theta(d)$ samples, and there exist polynomial time algorithms [CM20, TDD+25, DDM+25, KZM25, DLB25] that achieve this for generic generative-exponent-two multi-index models which covers almost all canonical examples. Nevertheless, in general, it is not clear whether neural networks can achieve this algorithmic limit and, if so, how. The earliest work on neural networks learning of multi-index models [DLS22] showed that standard two-layer neural networks can efficiently learn a large subclass of multi-index models with a sub-optimal $\widetilde{\Theta}(d^{2})$ sample size. Subsequently, [AAM22][AAM23] studied the problem of neural network learning multi-index models with a staircase structure. However, their results are far from optimal and are restricted to special classes of functions. The closest attempt to the general problem is [ADK+25], which shows that a two-layer network trained via gradient descent can weakly recover some of the features in a multi-index model using $\widetilde{\Theta}(d)$ samples. Nevertheless, since the features are only weakly recovered and are not guaranteed to cover the entire span, their results do not establish that these features can be directly leveraged to successfully learn the targets in later stages of training. Therefore, the following question remains open.
@@ 1.1 Main Contributions
We note that we use a standard initialization scheme for the neural network parameters, without using a spectral initialization or a hot start as in [MUZ21, CS25]. The neural networks can only learn the right features from and only from gradient descent, trained for a divering number of steps. This time scale requirement rule out all the approaches based on DMFT or state evolution, such as [ABUZ18, BES+22, CCM25, GTM+23, BMZ24] which requires either information/generative exponent one or a hot start/spectral initialization, to guarantee the learning timescale can be made constant. We also highlight that the power iteration approximation does not work if the learning rate is too small. It is important to choose a moderately large learning rate. For a more detailed discussion, please see Section 4.
@@ 1.2 Related Works
To start with, [BKM+19, MM18, LL19, LAL19, MKLZ21] studies generic single-index models and sharply derives the learning threshold. In particular, [MM18, LL19, LAL19, MKLZ21] show that for a broad class of single-index models, spectral methods with suitable data pre-processing can recover the hidden direction at the optimal threshold, matching the performance of an optimal approximate message passing (AMP) [BKM+19]. In contrast, [AGJ21] considers a more restrictive setting and examines the learning of single-index models when using online SGD as the optimizer without data preprocessing, and introduces the concept of an information exponent, which characterizes the learning complexity in their setup.
A substantial line of works studies how to extend these results to multi-index models such as [CM20, AAM22, AAM23, TDD+25, DLB25, KZM25, DDM+25]. On one hand, [ABAM24, ABAM23] introduces the notion of leap complexity, showing that hierarchical/staircase structure in multi-index models can facilitate learning. On the other hand, [DPVLB24, DLB25] extend this line of work by formalizing the generative exponent for single-index models and the generative leap exponent for multi-index models, extending the previous results to higher order and incorporating the benefits of data pre-processing. In particular, we highlight [DLB25] for its two key contributions.
@@
We note that these papers are in general not tight in terms of leading order constants.
Subsequently, for the precise (weak recovery) learning threshold for multi-index models, [TDD+25] non-rigorously derives the sharp optimal sample size required for weakly learning generative-leap-exponent-two multi-index models by performing a local analysis of Bayes AMP, extending the results of [BKM+19, MM18] for single-index models when the generative exponent of the target equals two. Moreover, as two follow-up works, [KZM25, DDM+25] study learning multi-index models using an optimal spectral method, although their sample thresholds in [KZM25] are in general not always optimal, and none of those works can deal with the generative-staircase setup. We emphasize that the above works depend on the special structure of multi-index models and therefore do not use general-purpose algorithms such as learning with standard neural networks trained via gradient descent.
@@
The term feature learning refers to the following phenomenon. In practice, neural networks trained with local gradient-based algorithms such as gradient descent can adapt to the low-dimensional structure of the tasks and efficiently learn useful hidden features, which in turn facilitates successful learning and enables effective transfer learning to downstream tasks. Among the first several works in this direction, as we have mentioned previously, [DLS22] demonstrates that neural networks can learn generic multi-index models at a sub-optimal sample size of $n=\widetilde{\Theta}(d^{2})$ under some additional assumptions. Similarly, [ABAM24, ABAM23] establish a sub-optimal rate of $n=\widetilde{{\mathcal{O}}}(d^{L})$ for neural network learning, where $L$ denotes the leap complexity. [BES+22] studies the problem under a much more restrictive assumption of an information exponent one, and in this setting shows that neural networks can succeed with ${\mathcal{O}}(d)$ samples.
@@
Furthermore, as noted earlier, [LOSW24] shows that neural networks can learn generic single-index models with generative exponent at most two with sample complexity $\widetilde{{\mathcal{O}}}(d)$ through data re-using while [MU25] provides a sharp DMFT theory analyzing the learning dynamics of single-index models using a standard two-layer neural network when the sample size and input dimension are proportional. Together, these two works essentially conclude the studies of learning single-index models with neural networks. Next, [DTA+24] demonstrates that neural networks can learn multi-index models with generative leap exponent one using $n={\mathcal{O}}(d)$ samples in constant time, leveraging DMFT theory. Finally, as noted previously, [ADK+25] provides a loose result, showing that neural networks with ${\mathcal{O}}(d\log^{2}d)$ samples can recover the hidden subspace of multi-index models but in a weak recovery sense when the generative leap exponent of the target equals two. Besides linear features, we note that there is also a substantial line of work on learning nonlinear features with deep neural networks, including [NDL23, RZG23, WNL23, FWNL24].
@@
We are aware of and highlight this recent concurrent work [MW25]. Our works differ in the following aspects. [MW25] studies learning a multi-index model using a two-layer network trained with standard gradient descent instead of modified gradient-based algorithms. Their arguments are performed in the proportional regime, that is, $n/d\rightarrow\delta\in(0,+\infty)$, leveraging toolboxes from discrete time DMFT and random matrix theory. Besides, their results are sharp in constants, and they derive the sharp phase transition threshold at which the trained neural network can start to weakly recover the first hard direction (with generative exponent two). In contrast, [MW25] does not establish full subspace recovery and therefore cannot provide an end-to-end analysis of the problem. Instead our work provides an end-to-end learning guarantee via a different mechanism, albeit at the cost of a slightly worse sample complexity and the use of a modified layer-wise training scheme.
@@ 2.1 Discussion of Assumptions
The central assumption in our analysis is Assumption 5. Informally, this assumption requires that the second-order information of the data remain non-degenerate after applying the pre-processing map $y\mapsto\ell^{\prime}(0,y)$. We admit that for one fixed loss function $\ell$, it is straightforward to construct a target task such that the second-order information of $\ell^{\prime}(0,y)$ becomes degenerate. In practice, however, the target task is typically fixed, and one has the flexibility to choose among different loss functions to learn it. Recent studies on generative exponents and learning thresholds [TDD+25, DPVLB24, KZM25, DLB25, DDM+25] show that for a large subclass of common tasks, including all polynomial targets with non-generative-staircase structure, there exists a pre-processing function that ensures the second-order information is non-degenerate, and importantly, such a function need not be highly specialized. Consequently, we expect that $\ell^{\prime}(0,y)$ exhibits non-degenerate second-order information for at least some of those commonly used loss functions in practice, and at least this already includes all targets in [DLS22]. We will expand on this point in Section 5.
@@ 2.2 Training Algorithm
We introduce our layer-wise gradient descent algorithm. At initialization, we set $b_{j}$ as zero and impose a symmetric structure on $a_{j}$ and ${\bm{w}}_{j}$. Specifically, let the width $m$ be even, and for all $j\in[m]$ we set
@@
The network is trained via layer-wise gradient descent with sample splitting, a common approach in theoretical analysis such as [WNL23, NDL25]. To start with, we generate two independent datasets $\mathcal{D}_{i}$ where $i=1,2$ and each consisting of $n$ independent samples. The training algorithm proceeds as follows. we first train $\bm{W}$} via gradient descent for $\mathrm{T}_{1}$ steps on the empirical loss $\widehat{\mathcal{L}}_{\mathcal{D}_{1}}({\bm{\Theta}})$, then reinitialize $\bm{b}$} and train $\bm{a}$ via gradient descent for $\mathrm{T}_{2}$ steps on the empirical loss $\widehat{\mathcal{L}}_{\mathcal{D}_{2}}({\bm{\Theta}})$. The pseudocode of this procedure is given below in Algorithm 1.
@@ 4.2 Comparison with the Existing Approaches
First of all, we note that our results cannot be recovered using a DMFT or state evolution framework, such as [ABUZ18, BES+22, CCM25, GTM+23, BMZ24] and many more. Essentially, in order to learn those generative-exponent two directions under the sample size $\widetilde{\Theta}(d)$, one needs a diverging time horizon to escape the initial saddle point if there is no hot start or explicit spectral initialization. Under the DMFT/state evolution propotional limit, the neurons shall not escape the initial stationary point.
@@ 5 Further Discussions
The generative exponent is important since [DLB25] proves that the complexity of efficiently learning a multi-index model is $\Theta(d^{k^{*}/2})$ via providing both the LDP lower bound and a matching spectral algorithm upper bound. Further, it can be shown that almost all common multi-index models have the generative leap exponent $k^{*}\leq 2$. For instance, it works when the link function is 1) all the polynomials, 2) piecewise linear functions, 3) intersection of halfspace. Therefore, we can always extract information from computing $\mathbb{E}\big[\mathcal{T}(y,{\bm{z}}_{S_{i}})\mathrm{He}_{2}({\bm{z}}_{S_{i}^{\perp}})\big]$ with a proper pre-processing function $\mathcal{T}$. In addition, Lemma F.14 of [DPVLB24] indicates that $\mathcal{T}$ need not to be highly specific. In fact, they prove that even random polynomials suffice in the single-index setting.
@@ 6.2 Proper Choice of Loss Function
7 Conclusions and Future Works
In this paper, we demonstrate that under Gaussian input, a two-layer neural network trained with layer-wise gradient descent can learn generic multi-index models up to $o_{d}(1)$ test error using $\widetilde{{\mathcal{O}}}(d)$ sample and $\widetilde{{\mathcal{O}}}(d^{2})$ time complexity, which matches the information-theoretic threshold up to the leading order. We emphasize that our analysis does not rely on any particular choice of activation, loss, or link function; rather, the assumptions imposed are generic and hold for a broad class of functions. A key and somewhat unexpected insight from the proof is that the inner layer must be trained for more than a constant number of steps to successfully learn the hidden features, in order to recover the hidden features most efficiently and thus achieve optimal performance. This observation is impossible to be observed via a DMFT or state evolution framework over bounded time horizon, and has the potential to extend to other settings. Overall, our results demonstrate that neural networks are capable of effective representation learning and can fully learn the target function on this canonical task with near-optimal efficiency.
Looking ahead, we believe that these insights may extend beyond the Gaussian setting to more general input distributions. Specifically, consider the target $f({\bm{x}})=g({\bm{U}}{\bm{x}})$ with $\bm{x}$} drawn from a generic distribution having finite second moments, and assume that $\bm{U}$} satisfies a suitable incoherence condition. Intuitively, even if $\bm{x}$} is not Gaussian, the projection ${\bm{U}}{\bm{x}}$ should be approximately Gaussian due to the central limit theorem, with incoherence providing the necessary averaging effect. We conjecture that our results remain valid in this broader setting, which presents an intriguing direction for future work.

### v2

@@ 1 Introduction
Modern large language models powered by deep neural networks [KSH12, HZR+16, VSP+23] have shown strong empirical success across many domains. It is believed that much of the effectiveness of the pretraining–finetuning paradigm comes from their ability to extract useful representations from data. In practice, neural networks often first capture important hidden features, which helps them to learn complex tasks more efficiently. This adaptability offers a clear performance advantage over conventional fixed feature methods, such as kernel techniques [WLL+20, AL20, BL20].
@@
Despite these empirical successes, the theoretical understanding of why neural networks can efficiently and adaptively learn important features from data remains largely incomplete. Significant effort has been devoted in this direction, with a series of works [DLS22, BES+22, BMZ24, LOS+24, MU25, MW26] investigating how neural networks can uncover hidden features and subsequently learn single-index and multi-index models, where the labels depend on the $d$-dimensional data inputs only through their projection onto a one-dimensional or low-dimensional subspace, respectively. The study of learning single-index models via neural networks is relatively well developed. Notably, [LOS+24] show that neural networks can successfully learn single-index models with $\widetilde{\mathcal{O}}(d)$ sample complexity, which is optimal up to leading order, while [MU25] provides a sharp DMFT theory analyzing the learning dynamics of single-index models using a two-layer neural network when the sample size and input dimension are proportional.
@@
However, a single-index model involves only one hidden feature, whereas in practice we often expect multiple distinct features, motivating the study of neural networks learning of multi-index models. On the contrary, positive results on neural networks learning of multi-index models remain rather limited. Note first that the information-theoretic limit for learning multi-index models is still $\Theta(d)$ samples, and there exist polynomial time algorithms [CM20, TDD+25, DDM+25, kovačević2025spectralestimatorsmultiindexmodels, DLB25] that achieve this for generic generative-exponent-two multi-index models which covers almost all canonical examples. Nevertheless, in general, it is not clear whether neural networks can achieve this algorithmic limit and, if so, how. The earliest work on neural networks learning of multi-index models [DLS22] showed that standard two-layer neural networks can efficiently learn a large subclass of multi-index models with a sub-optimal $\widetilde{\Theta}(d^{2})$ sample size. Subsequently, [AAM22][AAM23a] studied the problem of neural network learning multi-index models with a staircase structure. However, their results are far from optimal and are restricted to special classes of functions. The closest attempt to the general problem is [MW26], which shows that a two-layer network trained via gradient descent can weakly recover some of the features in a multi-index model using $\Theta(d)$ samples. Nevertheless, since the features are only weakly recovered and are not guaranteed to cover the entire span, their results do not establish that these features can be directly leveraged to successfully learn the targets in later stages of training. Therefore, the following question remains open.
@@ 1.1 Main contributions
We note that we use a standard initialization scheme for the neural network parameters, without using a spectral initialization or a hot start as in [MUZ21, CS25]. The neural networks can only learn the right features from and only from gradient descent, trained for a divering number of steps. This time scale requirement rule out all the approaches based on DMFT or state evolution, such as [ABU+18, BES+22, CCM25, GTM+23, BMZ24] which requires either information/generative exponent one or a hot start/spectral initialization, to guarantee the learning timescale can be made constant. We also highlight that the power iteration approximation does not work if the learning rate is too small. It is important to choose a moderately large learning rate. For a more detailed discussion, please see Section 4.
@@ 1.2 Related works
To start with, [BKM+19, MM18, LL19, LAL19, MKL+21] studies generic single-index models and sharply derives the learning threshold. In particular, [MM18, LL19, LAL19, MKL+21] show that for a broad class of single-index models, spectral methods with suitable data pre-processing can recover the hidden direction at the optimal threshold, matching the performance of an optimal approximate message passing (AMP) [BKM+19]. In contrast, [AGJ21] considers a more restrictive setting and examines the learning of single-index models when using online SGD as the optimizer without data preprocessing, and introduces the concept of an information exponent, which characterizes the learning complexity in their setup.
A substantial line of works studies how to extend these results to multi-index models such as [CM20, AAM22, AAM23a, TDD+25, DLB25, kovačević2025spectralestimatorsmultiindexmodels, DDM+25, DIK+25]. On one hand, [ABM24, ABM23b] introduces the notion of leap complexity, showing that hierarchical/staircase structure in multi-index models can facilitate learning. On the other hand, [DPL+24, DLB25] extend this line of work by formalizing the generative exponent for single-index models and the generative leap exponent for multi-index models, extending the previous results to higher order and incorporating the benefits of data pre-processing. In particular, we highlight [DLB25] for its two key contributions.
@@
These papers are in general not tight in terms of leading order constants. We also note that [DIK+25] introduces a related notion, i.e., the $m$-well-behaved multi-index model, which coincides with the generative exponent in many cases. While their sample complexity exhibits a worse dependence on $d$ compared to [DLB25], their work offers two advantages: it targets agnostic PAC learning rather than subspace recovery, and it provides uniform sample complexity guarantees that apply agnostically to all multi-index models within a certain given class.
Subsequently, for the precise (weak recovery) learning threshold for multi-index models, [TDD+25] non-rigorously derives the sharp optimal sample size required for weakly learning generative-leap-exponent-two multi-index models by performing a local analysis of Bayes AMP, extending the results of [BKM+19, MM18] for single-index models when the generative exponent of the target equals two. Moreover, as two follow-up works, [kovačević2025spectralestimatorsmultiindexmodels, DDM+25] study learning multi-index models using an optimal spectral method, although their sample thresholds in [kovačević2025spectralestimatorsmultiindexmodels] are in general not always optimal, and none of those works can deal with the generative-staircase setup. We emphasize that the above works depend on the special structure of multi-index models and therefore do not use general-purpose algorithms such as learning with standard neural networks trained via gradient descent.
@@
The term feature learning refers to the following phenomenon. In practice, neural networks trained with local gradient-based algorithms such as gradient descent can adapt to the low-dimensional structure of the tasks and efficiently learn useful hidden features, which in turn facilitates successful learning and enables effective transfer learning to downstream tasks. Among the first several works in this direction, as we have mentioned previously, [DLS22] demonstrates that neural networks can learn generic multi-index models at a sub-optimal sample size of $n=\widetilde{\Theta}(d^{2})$ under some additional assumptions. Similarly, [ABM24, ABM23b] establish a sub-optimal rate of $n=\widetilde{{\mathcal{O}}}(d^{L})$ for neural network learning, where $L$ denotes the leap complexity. [BES+22] studies the problem under a much more restrictive assumption of an information exponent one, and in this setting shows that neural networks can succeed with ${\mathcal{O}}(d)$ samples.
@@
Furthermore, as noted earlier, [LOS+24] shows that neural networks can learn generic single-index models with generative exponent at most two with sample complexity $\widetilde{{\mathcal{O}}}(d)$ through data re-using while [MU25] provides a sharp DMFT theory analyzing the learning dynamics of single-index models using a standard two-layer neural network when the sample size and input dimension are proportional. Together, these two works essentially conclude the studies of learning single-index models with neural networks. Next, [DTA+24] demonstrates that neural networks can learn multi-index models with generative leap exponent one using $n={\mathcal{O}}(d)$ samples in constant time, leveraging DMFT theory. Finally, as noted previously, [ADK+25] provides a loose result, showing that neural networks with ${\mathcal{O}}(d\log^{2}d)$ samples can recover the hidden subspace of multi-index models but in a weak recovery sense when the generative leap exponent of the target equals two. Besides linear features, we note that there is also a substantial line of work on learning nonlinear features with deep neural networks, including [NDL23, RZG23, WNL23, FWN+24].
@@
We are aware of and highlight this recent concurrent work [MW26]. Our works differ in the following aspects. [MW26] studies learning a multi-index model using a two-layer network trained with standard gradient descent instead of modified gradient-based algorithms. Their arguments are performed in the proportional regime, that is, $n/d\rightarrow\delta\in(0,+\infty)$, leveraging toolboxes from discrete time DMFT and random matrix theory. Besides, their results are sharp in constants, and they derive the sharp phase transition threshold at which the trained neural network can start to weakly recover the first hard direction (with generative exponent two). In contrast, [MW26] does not establish full subspace recovery and therefore cannot provide an end-to-end analysis of the problem. Instead our work provides an end-to-end learning guarantee via a somehow different mechanism, albeit at the cost of a slightly worse sample complexity and the use of a modified layer-wise training scheme.
@@ 2.1 Discussion of assumptions
The central assumption in our analysis is Assumption 5. Informally, this assumption requires that the second-order information of the data remain non-degenerate after applying the pre-processing map $y\mapsto\ell^{\prime}(0,y)$. We admit that for one fixed loss function $\ell$, it is straightforward to construct a target task such that the second-order information of $\ell^{\prime}(0,y)$ becomes degenerate. In practice, however, the target task is typically fixed, and one has the flexibility to choose among different loss functions to learn it. Recent studies on generative exponents and learning thresholds [TDD+25, DPL+24, kovačević2025spectralestimatorsmultiindexmodels, DLB25, DDM+25] show that for a large subclass of common tasks, including all polynomial targets with non-generative-staircase structure, there exists a pre-processing function that ensures the second-order information is non-degenerate, and importantly, such a function need not be highly specialized. Consequently, we expect that $\ell^{\prime}(0,y)$ exhibits non-degenerate second-order information for at least some of those commonly used loss functions in practice, and at least this already includes all targets in [DLS22]. We will expand on this in Section 5.
@@ 2.2 Training algorithm
We introduce our layer-wise gradient descent algorithm.
At initialization, we set $b_{j}$ as zero and impose a symmetric structure on $a_{j}$ and ${\bm{w}}_{j}$. Specifically, let the width $m$ be even, and for all $j\in[m]$ we set
@@
The network is trained via layer-wise gradient descent with sample splitting, a common approach in theoretical analysis such as [WNL23, NDL25]. To start with, we generate two independent datasets $\mathcal{D}_{i}$ where $i=1,2$ and each consisting of $n$ independent samples. The training algorithm proceeds as follows. we first train $\bm{W}$} via gradient descent for $\mathrm{T}_{1}$ steps on the empirical loss $\widehat{\mathcal{L}}_{\mathcal{D}_{1}}({\bm{\Theta}})$, then reinitialize $\bm{b}$} and train $\bm{a}$ via gradient descent for $\mathrm{T}_{2}$ steps on the empirical loss $\widehat{\mathcal{L}}_{\mathcal{D}_{2}}({\bm{\Theta}})$. The pseudocode of this procedure is given below.
@@ 4.2 Comparison with the existing approaches
First of all, we note that our results cannot be recovered using a DMFT or state evolution framework, such as [ABU+18, BES+22, CCM25, GTM+23, BMZ24] and many more. Essentially, in order to learn those generative-exponent two directions under the sample size $\widetilde{\Theta}(d)$, one needs a diverging time horizon to escape the initial saddle point if there is no hot start or explicit spectral initialization. Under the DMFT/state evolution propotional limit, the neurons shall not escape the initial stationary point.
@@ 5 Further discussions
The generative exponent is important since [DLB25] proves that the complexity of efficiently learning a multi-index model is $\Theta(d^{k^{*}/2})$ via providing both the LDP lower bound and a matching spectral algorithm upper bound. Further, it can be shown that almost all common multi-index models have the generative leap exponent $k^{*}\leq 2$. For instance, it works when the link function is 1) all the polynomials, 2) piecewise linear functions, 3) intersection of halfspace. Therefore, we can always extract information from computing $\mathbb{E}\big[\mathcal{T}(y,{\bm{z}}_{S_{i}})\mathrm{He}_{2}({\bm{z}}_{S_{i}^{\perp}})\big]$ with a proper pre-processing function $\mathcal{T}$. In addition, Lemma F.14 of [DPL+24] indicates that $\mathcal{T}$ need not to be highly specific. In fact, they prove that even random polynomials suffice in the single-index setting.
@@ 6.2 Proper choice of loss function
7 Conclusions
In this paper, we demonstrate that under Gaussian input, a two-layer neural network trained with layer-wise gradient descent can learn generic multi-index models up to $o_{d}(1)$ test error using $\widetilde{{\mathcal{O}}}(d)$ sample and $\widetilde{{\mathcal{O}}}(d^{2})$ time complexity, which matches the information-theoretic threshold up to the leading order. We emphasize that our analysis does not rely on any particular choice of activation, loss, or link function; rather, the assumptions imposed are generic and hold for a broad class of functions. An important and somewhat unexpected insight from the proof is that the inner layer must be trained for more than a constant number of steps to successfully learn the hidden features, in order to recover the hidden features most efficiently and thus achieve optimal performance. This observation is impossible to be observed via a DMFT or state evolution framework over bounded time horizon, and has the potential to extend to other settings. Overall, our results demonstrate that neural networks are capable of effective representation learning and can fully learn the target function on this canonical task with near-optimal efficiency.
We believe that these insights may extend beyond the Gaussian setting to more general input distributions. Specifically, consider the target $f({\bm{x}})=g({\bm{U}}{\bm{x}})$ with $\bm{x}$} drawn from a generic distribution having finite second moments, and assume that $\bm{U}$} satisfies a suitable incoherence condition. Intuitively, even if $\bm{x}$} is not Gaussian, the projection ${\bm{U}}{\bm{x}}$ should be approximately Gaussian due to the central limit theorem, with incoherence providing the necessary averaging effect. We conjecture that our results remain valid in this broader setting, which presents an intriguing direction for future work.